# TaoSearchEmb: A Multi-Objective Reinforcement Learning Framework for Dense Retrieval in Taobao Search

### v1

@@ Abstract.
Dense retrieval, as the core component of e-commerce search engines, maps user queries and items into a unified semantic space through pre-trained embedding models to enable large-scale real-time semantic retrieval. Despite the rapid advancement of LLMs gradually replacing traditional BERT architectures for embedding, their training paradigms still adhere to BERT-like supervised fine-tuning and hard negative mining strategies. This approach relies on complex offline hard negative sample construction pipelines, which constrain model iteration efficiency and hinder the evolutionary potential of semantic representation capabilities. Besides, existing multi-task learning frameworks face the seesaw effect when simultaneously optimizing semantic relevance and non-relevance objectives. In this paper, we propose Retrieval-GRPO, a multi-objective reinforcement learning-based dense retrieval framework designed to address these challenges. The method eliminates offline hard negative sample construction by dynamically retrieving Top-K candidate products for each query during training, while introducing a relevance LLM as a reward model to generate real-time feedback. Specifically, the retrieval model dynamically optimizes embedding representations through reinforcement learning, with reward signals combining LLM-generated relevance scores, product quality scores, and multi-way exclusivity metrics to achieve multi-objective user preference alignment and real-time error correction. This mechanism not only removes dependency on hard negatives but also mitigates the seesaw effect through collaborative multi-objective optimization, significantly enhancing the model’s semantic generalization capability for complex long-tail queries. Extensive offline and online experiments validate the effectiveness of Retrieval-GRPO, which has been deployed on China’s largest e-commerce platform.
@@ 1. Introduction
Dense retrieval(Guo_Fan_Ai_Croft_2016)(Guo_Fan_Pang_Yang_Ai_Zamani_Wu_Croft_Cheng_2020)(Mitra_Craswell_2017) has emerged as a critical technology in e-commerce search engines to address the unique challenge of precisely aligning massive items with user queries. This approach employs advanced pre-trained embedding models to encode user queries and item descriptions into a unified semantic vector space, enabling real-time semantic search over billions of items via Approximate Nearest Neighbor (ANN) algorithms(malkov2018efficient). The methodology effectively learns query-item alignment relationships within a unified semantic space, even in the absence of direct lexical overlap between them. Dense retrieval directly improves users’ search experience and merchant exposure efficiency, thereby driving significant improvements in overall platform transaction conversion rates.
@@
Previous approaches primarily relied on encoder-only architectures like BERT(Devlin_Chang_Lee_Toutanova_2019) and RoBERTa(Liu_Ott_Goyal_Du_Joshi_Chen_Levy_Lewis_Zettlemoyer_Stoyanov_2019), utilizing the bidirectional attention mechanism to produce contextual semantic embeddings. Benefiting from rapid advancements in large language model technologies, emerging approaches such as Gemini Embedding(lee2025gemini), Seed Embedding(guo2025seed1), QWEN3-Embedding(zhang2025qwen3), and NV-Embed(lee2024nv) employ LLMs as encoders to generate semantic embeddings for retrieval, leveraging their extensive knowledge bases and zero-shot generalization capabilities. Despite incorporating LLMs, most current methods still adopt training paradigms designed for BERT-like models, enhancing performance on long-tail challenging queries through continued supervised fine-tuning (SFT) and persistent hard negative sample mining. Notably, offline hard negative sample mining constitutes a complex systemic workflow, which restricts iteration speed and limits the evolution potential of representation capabilities.
@@
Furthermore, in e-commerce search systems, semantic relevance between queries and items is not the sole evaluation criterion. Existing methods typically employ multi-task learning frameworks when modeling non-relevance objectives. However, the simultaneous optimization of multiple losses suffers from the seesaw phenomenon(10.1145/3383313.3412236), which constrains the model’s performance ceiling.
@@
In this paper, we propose Retrieval-GRPO, a multi-objective reinforcement learning framework for dense retrieval. This approach eliminates the need for offline manual construction of hard negative samples. Instead, during training, we dynamically retrieve top-k items for each query and employ a relevance LLM(dong2025taosr1) as the reward model to generate scores. These relevance scores are combined with item quality scores and exclusivity metrics to form multi-objective rewards, enabling real-time user preference alignment and error correction through RL training.
@@
In summary, compared to conventional dense retrieval paradigms, R-GRPO offers three key contributions:
@@ 2.1. Dense Retrieval
Dense retrieval encodes queries and items into a shared semantic space and retrieves the top-N most similar items using approximate nearest neighbor (ANN) algorithms(malkov2018efficient). Early dense retrieval methods primarily relied on encoder-only pre-trained models like BERT(Devlin_Chang_Lee_Toutanova_2019), RoBERTa(Liu_Ott_Goyal_Du_Joshi_Chen_Levy_Lewis_Zettlemoyer_Stoyanov_2019). SimCSE(gao2021simcse) first introduced a contrastive learning framework for dense retrieval tasks, using dropout as an unsupervised self-prediction technique in its contrastive leaning framework, after that the contrastive learning paradigm became widely adopted. Beyond pure text representation, CLIP(radford2021learning) leverages a large-scale image-text matching data to pretrain multi-modal representations through simple contrastive learning objectives for multimodal alignment tasks. With the rapid development of large language models, many studies now employ LLMs as backbones for dense retrieval, leveraging their extensive world knowledge and zero-shot generalization capabilities to obtain superior universal semantic representations. Notably, RepLLaMA(ma2024fine) first demonstrated the effectiveness of multi-stage supervised fine-tuning (SFT) of large language models for text retrieval. NV-Embed(lee2024nv) proposes a latent attention layer to obtain pooled embeddings and enhances model performance through contrastive learning with in-batch negatives and curated hard negative examples. Gemini Embedding(lee2025gemini) utilizes Gemini’s multilingual and code comprehension capabilities to generate highly versatile embeddings across languages and text modalities. Seed Embedding(guo2025seed1) adopts a Mixture-of-Experts (MoE) model as its backbone, supporting multiple embedding dimensions via Matryoshka Representation Learning (MRL)(kusupati2022matryoshka) training similar to NV-Embed, and employs two-stage contrastive learning to strengthen universal representation capabilities. Gte-Qwen(li2023towards) integrates bidirectional attention mechanisms and enhances performance through instruction tuning using both weakly supervised and supervised data. Building on gte-Qwen, Qwen3 Embedding(zhang2025qwen3) introduces an innovative multi-stage training pipeline that combines large-scale unsupervised pre-training with supervised fine-tuning on high-quality datasets, ensuring robustness and adaptability through effective model merging strategies. Current LLM-based dense retrieval models achieve state-of-the-art (SOTA) performance on the MTEB leaderboard(muennighoff2023mteb), demonstrating the effectiveness of large models for semantic representation.
Figure 1. The Retrieval-GRPO overall training framework consists of 3 steps: (1)Candidates Selection; (2)Multi-Objective Reward Calculation; (3)GRPO Optimation:
@@ 2.2. Reinforcement Learning for LLM
In recent years, large language model (LLM) technology has rapidly advanced, with reinforcement learning (RL)(wang2024reinforcement) being widely integrated into LLM training pipelines as a key technical approach to enhance model performance and controllability. The core application of RL in LLMs involves modeling text generation as a sequential decision-making problem, where the model acts as an agent, generated text represents actions, and external feedback forms reward signals. Among these approaches, Reinforcement Learning from Human Feedback (RLHF) has quickly become the dominant paradigm. OpenAI’s Proximal Policy Optimization (PPO)(Schulman_Wolski_Dhariwal_Radford_Klimov_2017) stands as one of the most representative algorithms in the RLHF framework. Subsequently, OpenAI first applied PPO to align human preferences in its InstructGPT(ouyang2022training). In PPO, the policy model serves as the optimization target, generating sampled text from input prompts. The reward model, trained on human preference data, provides scalar reward signals based on response quality from the policy model. A reference model calculates KL divergence penalties to prevent excessive deviation from the original model, while a value model estimates action value functions representing expected cumulative future value. The PPO algorithm coordinates these four components to align human preferences. Despite PPO’s success, its complex workflow and high computational costs persist. The Direct Preference Optimization (DPO)(rafailov2023direct) algorithm addresses this by assuming human preferences follow the Bradley-Terry model, simplifying preference learning into a directly optimizable classification problem through maximum likelihood objectives. While DPO achieves stable performance with higher efficiency and outperforms PPO in some tasks, PPO maintains advantages in complex multi-step reasoning tasks. To reduce PPO’s complexity and computational demands, DeepSeek proposes the GRPO(shao2024deepseekmath) algorithm during the training of its DeepSeekMath. GRPO eliminates the value model in PPO and adopts a group-sampling-based baseline estimation method, using average group rewards as the baseline to calculate inter-group advantages, significantly reducing computational overhead. Building on PPO, LitePPO(liu2025part) discovered that simply applying robust advantage normalization and token-level loss aggregation techniques to vanilla PPO could consistently improve performance. Currently, the RL research continues to drive progress in LLM training, providing sustainable iterative methods for aligning large language models with human preferences.
@@ 3. Methods
In modern e-commerce search systems, dense retrieval aims to map queries and items into a shared semantic embedding space, enabling efficient selection through geometric proximity. Given a query $q_{i}$ and an item collection $\mathcal{D}=\{d_{1},\ldots,d_{n}\}$, the framework typically employs a dual-encoder architecture(reimers2019sentence) to encode $q_{i}$ and $\mathcal{D}$ to a semantic embedding as Equation 1 and 2.
@@
In this section, we detail our multi-objective dense retrieval training framework, which includes Supervised Fine-Tuning (SFT) and Retrieval-GRPO. Additionally, we provide a detailed separate introduction to the Reward design in Retrieval-GRPO. First, SFT is performed to endow the model with discriminative capabilities on simple positive and negative samples. Then, through Retrieval-GRPO, the model learns from the reward of multi-objective user preferences for real-time correction, eliminating the need for traditional manual construction of hard negative samples(li2024syneg).
@@ 3.1. SFT
Traditionally, in the SFT phase of dense retrieval, training samples consist of ¡query, item¿ pairs for positive samples, while negative samples are not specifically collected but obtained through in-batch negative sampling(radford2021learning). The InfoNCE loss is used to pull closer the distance between positive query-item pairs and push away irrelevant query-item pairs in the semantic embedding space. The InfoNCE loss is as :
@@
However, this approach leads to only those items present as positive samples in the training set being utilized as negative samples for other queries. Meanwhile, items in the product pool that lack historical user interactions are entirely ignored during training. To address this, we enhance the conventional InfoNCE in-batch negative sampling by introducing global negative sampling, which randomly selects items from the product pool to serve as negative samples for queries. Consequently, the negative samples are derived partially from in-batch sampling and partially from the product pool, ensuring that all items in the product pool are exposed to the model during training. The loss is as :
@@ 3.2. Retrieval-GRPO
  • Candidate Selection: The dense retrieval model encodes queries and items into a unified semantic embedding space. Using the Approximate Nearest Neighbor (ANN) algorithm(malkov2018efficient), it retrieves the top-k items with the closest distances as candidate items.
@@ 3.2.1. Multi-Objective Reward
To satisfy users’ search experience, dense retrieval aim to prioritize items that are highly relevant to the user’s query and of superior quality. However, due to the prevalence of multi-way retrieval architectures in e-commerce search systems, we particularly emphasize incrementally exclusive relevant items—those that are semantically related to the query but not exact matches—that provide incremental value compared to results from other retrieval ways.
@@
where the $f_{\text{relevance}}(q,d)$ denotes the relevance score between query and item computed by TaoSR1, $g_{\text{quality}}(d)$ denotes represents the objective quality score of items derived from historical transaction data, and $h_{\text{exclusivity}}(q,d)$ measures exclusivity against other retrieval-ways, indicating whether the item provides incremental value by being retrieved through the current channel but not through alternative retrieval ways.
  • Relevance reward:The query-item relevance score is computed by TaoSR1(dong2025taosr1), an advanced internal downstream relevance model deployed independently. TaoSR1 is a 42B-MoE LLM trained through multi-stage RL learning. We leverage it as the reward model in our reinforcement learning framework to provide real-time relevance scoring for query-item candidates generated by dense retrieval model.
@@ 3.2.2. Candidates Selection
In the Retrieval-GRPO framework, instead of offline hard negative samples construction, we treat the top-k items retrieved by the dense retrieval model as candidate predictions requiring correction. Leveraging the relevance reward model and multi-objective reward function described in Section 3.2.1, the rewards is computed in real time. Items within the top-k that receive low reward scores are penalized to dynamically adjust the model’s predictions.
@@ 3.2.3. GRPO loss
(7) $\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\bigg\{\min\bigg[\frac{\pi_{\theta}(s(q,d_{i})|q,d_{i})}{\pi_{\theta_{\text{old}}}(s(q,d_{i})|q,d_{i})}\hat{A}_{i,t},\\ &\quad\quad\text{clip}\left(\frac{\pi_{\theta}(s(q,d_{i})|q,d_{i})}{\pi_{\theta_{\text{old}}}(s(q,d_{i})|q,d_{i})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\bigg]-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}||\pi_{\text{ref}}\big]\bigg\}\Bigg]\end{split}$

### v2

# Retrieval-GRPO: A Multi-Objective Reinforcement Learning Framework for Dense Retrieval in Taobao Search

@@ Abstract.
Dense retrieval, as the core component of e-commerce search engines, maps user queries and items into a unified semantic space through pre-trained embedding models to enable large-scale, real-time semantic retrieval. Despite the rapid advancement of large language models (LLMs) gradually replacing traditional BERT architectures for embedding, their training paradigms still adhere to BERT-like supervised fine-tuning and hard negative mining strategies. This approach relies on complex offline hard negative sample construction pipelines, which constrain model iteration efficiency and hinder the evolutionary potential of semantic representation capabilities. Furthermore, existing multi-task learning frameworks face the seesaw effect when simultaneously optimizing semantic relevance and non-relevance objectives, limiting model performance ceilings. In this paper, we propose Retrieval-GRPO, a multi-objective reinforcement learning-based dense retrieval framework designed to address these challenges. The method eliminates offline hard negative sample construction by dynamically retrieving Top-K candidate products for each query during training, while introducing a relevance LLM as a reward model to generate real-time feedback. Specifically, the retrieval model dynamically optimizes embedding representations through reinforcement learning, with reward signals combining LLM-generated relevance scores, product quality scores, and multi-way exclusivity metrics to achieve multi-objective user preference alignment and real-time error correction. This mechanism not only removes dependency on hard negatives but also mitigates the seesaw effect through collaborative multi-objective optimization, significantly enhancing the model’s semantic generalization capability for complex long-tail queries. Extensive offline and online experiments validate the effectiveness of Retrieval-GRPO, which has been deployed on China’s largest e-commerce platform.
@@ 1. Introduction
Dense retrieval(Guo et al., 2016)(Guo et al., 2020)(Mitra and Craswell, 2017) has emerged as a critical technology in e-commerce search engines to address the unique challenge of precisely aligning massive items with user queries. This approach employs advanced pre-trained embedding models to encode user queries and item descriptions into a unified semantic vector space, enabling real-time semantic search over billions of items via Approximate Nearest Neighbor (ANN) algorithms(Malkov and Yashunin, 2018). The methodology effectively learns query-item alignment relationships within a unified semantic space, even in the absence of direct lexical overlap between them. Dense retrieval directly improves users’ search experience and merchant exposure efficiency, thereby driving significant improvements in overall platform transaction conversion rates.
@@
Previous approaches primarily relied on encoder-only architectures like BERT(Devlin et al., 2019) and RoBERTa(Liu et al., 2019), utilizing the bidirectional attention mechanism to produce contextual semantic embeddings. Benefiting from rapid advancements in large language model technologies, emerging approaches such as Gemini Embedding(Lee et al., 2025), Seed Embedding(Guo et al., 2025), QWEN3-Embedding(Zhang et al., 2025), and NV-Embed(Lee et al., 2024) employ LLMs as encoders to generate semantic embeddings for retrieval, leveraging their extensive knowledge bases and zero-shot generalization capabilities. Despite incorporating LLMs, most current methods still adopt training paradigms designed for BERT-like models, enhancing performance on long-tail challenging queries through continued supervised fine-tuning (SFT) and persistent hard negative sample mining. Notably, offline hard negative sample mining constitutes a complex systemic workflow, which restricts iteration speed and limits the evolution potential of representation capabilities.
@@
Furthermore, in e-commerce search systems, semantic relevance between queries and items is not the sole evaluation criterion. Existing methods typically employ multi-task learning frameworks when modeling non-relevance objectives. However, the simultaneous optimization of multiple losses suffers from the seesaw phenomenon(Tang et al., 2020), which constrains the model’s performance ceiling.
@@
In this paper, we propose Retrieval-GRPO, a multi-objective reinforcement learning framework for dense retrieval. This approach eliminates the need for offline manual construction of hard negative samples. Instead, during training, we dynamically retrieve top-k items for each query and employ a relevance LLM(Dong et al., 2025) as the reward model to generate scores. These relevance scores are combined with item quality scores and exclusivity metrics to form multi-objective rewards, enabling real-time user preference alignment and error correction through RL training.
@@
In summary, compared to conventional dense retrieval paradigms, Retrieval-GRPO offers three key contributions:
@@ 2.1. Dense Retrieval
Dense retrieval encodes queries and items into a shared semantic space and retrieves the top-N most similar items using approximate nearest neighbor (ANN) algorithms(Malkov and Yashunin, 2018). Early dense retrieval methods primarily relied on encoder-only pre-trained models like BERT(Devlin et al., 2019), RoBERTa(Liu et al., 2019). SimCSE(Gao et al., 2021) first introduced a contrastive learning framework for dense retrieval tasks, using dropout as an unsupervised self-prediction technique in its contrastive learning framework. After that, the contrastive learning paradigm became widely adopted. Beyond pure text representation, CLIP(Radford et al., 2021) leverages a large-scale image-text matching dataset to pretrain multimodal representations through simple contrastive learning objectives for multimodal alignment tasks. With the rapid development of large language models, many studies now employ LLMs as backbones for dense retrieval, leveraging their extensive world knowledge and zero-shot generalization capabilities to obtain superior universal semantic representations. Notably, RepLLaMA(Ma et al., 2024) first demonstrated the effectiveness of multi-stage supervised fine-tuning (SFT) of large language models for text retrieval. NV-Embed(Lee et al., 2024) proposes a latent attention layer to obtain pooled embeddings and enhances model performance through contrastive learning with in-batch negatives and curated hard negative examples. Gemini Embedding(Lee et al., 2025) utilizes Gemini’s multilingual and code comprehension capabilities to generate highly versatile embeddings across languages and text modalities. Seed Embedding(Guo et al., 2025) adopts a Mixture-of-Experts (MoE) model as its backbone, supporting multiple embedding dimensions via Matryoshka Representation Learning (MRL)(Kusupati et al., 2022) training similar to NV-Embed, and employs two-stage contrastive learning to strengthen universal representation capabilities. The gte-Qwen(Li et al., 2023) integrates bidirectional attention mechanisms and enhances performance through instruction tuning using both weakly supervised and supervised data. Building on gte-Qwen, Qwen3 Embedding(Zhang et al., 2025) introduces an innovative multi-stage training pipeline that combines large-scale unsupervised pre-training with supervised fine-tuning on high-quality datasets, ensuring robustness and adaptability through effective model merging strategies. Current LLM-based dense retrieval models achieve state-of-the-art (SOTA) performance on the MTEB leaderboard(Muennighoff et al., 2023), demonstrating the effectiveness of large models for semantic representation.
Figure 1. The Retrieval-GRPO overall training framework consists of 3 steps: (1)Candidates Selection; (2)Multi-Objective Reward Calculation; (3)GRPO Optimization
@@ 2.2. Reinforcement Learning for LLM
In recent years, large language model (LLM) technology has rapidly advanced, with reinforcement learning (RL)(Wang et al., 2024) being widely integrated into LLM training pipelines as a key technical approach to enhance model performance and controllability. The core application of RL in LLMs involves modeling text generation as a sequential decision-making problem, where the model acts as an agent, generated text represents actions, and external feedback forms reward signals. Among these approaches, Reinforcement Learning from Human Feedback (RLHF) has quickly become the dominant paradigm. OpenAI’s Proximal Policy Optimization (PPO)(Schulman et al., 2017) stands as one of the most representative algorithms in the RLHF framework. Subsequently, OpenAI first applied PPO to align human preferences in its InstructGPT(Ouyang et al., 2022). In PPO, the policy model serves as the optimization target, generating sampled text from input prompts. The reward model, trained on human preference data, provides scalar reward signals based on response quality from the policy model. A reference model calculates KL divergence penalties to prevent excessive deviation from the original model, while a value model estimates action value functions representing expected cumulative future value. The PPO algorithm coordinates these four components to align human preferences. Despite PPO’s success, its complex workflow and high computational costs persist. The Direct Preference Optimization (DPO)(Rafailov et al., 2023) algorithm addresses this by assuming human preferences follow the Bradley-Terry model, simplifying preference learning into a directly optimizable classification problem through maximum likelihood objectives. While DPO achieves stable performance with higher efficiency and outperforms PPO in some tasks, PPO maintains advantages in complex multi-step reasoning tasks. To reduce PPO’s complexity and computational demands, DeepSeek proposes the GRPO(Shao et al., 2024) algorithm during the training of its DeepSeekMath. GRPO eliminates the value model in PPO and adopts a group-sampling-based baseline estimation method, using average group rewards as the baseline to calculate inter-group advantages, significantly reducing computational overhead. Building on PPO, LitePPO(Liu et al., 2025) discovered that simply applying robust advantage normalization and token-level loss aggregation techniques to vanilla PPO could consistently improve performance. Currently, the RL research continues to drive progress in LLM training, providing sustainable iterative methods for aligning large language models with human preferences.
@@ 3. Methods
In modern e-commerce search systems, dense retrieval aims to map queries and items into a shared semantic embedding space, enabling efficient selection through geometric proximity. Given a query $q_{i}$ and an item collection $\mathcal{D}=\{d_{1},\ldots,d_{n}\}$, the framework typically employs a dual-encoder architecture(Reimers and Gurevych, 2019) to encode $q_{i}$ and $\mathcal{D}$ to a semantic embedding as Equation 1 and 2.
@@
In this section, we detail our multi-objective dense retrieval training framework, which includes Supervised Fine-Tuning (SFT) and Retrieval-GRPO. Additionally, we provide a detailed, separate introduction to the Reward design in Retrieval-GRPO. First, SFT is performed to endow the model with discriminative capabilities on simple positive and negative samples. Then, through Retrieval-GRPO, the model learns from the reward of multi-objective user preferences for real-time correction, eliminating the need for traditional manual construction of hard negative samples(Li et al., 2024).
@@ 3.1. SFT
Traditionally, in the SFT phase of dense retrieval, training samples consist of ¡query, item¿ pairs for positive samples, while negative samples are not specifically collected but obtained through in-batch negative sampling(Radford et al., 2021). The InfoNCE loss is used to pull closer the distance between positive query-item pairs and push away irrelevant query-item pairs in the semantic embedding space. The InfoNCE loss is as follows:
@@
However, this approach leads to only those items present as positive samples in the training set being utilized as negative samples for other queries. Meanwhile, items in the product pool that lack historical user interactions are entirely ignored during training. To address this, we enhance the conventional InfoNCE in-batch negative sampling by introducing global negative sampling, which randomly selects items from the product pool to serve as negative samples for queries. Consequently, the negative samples are derived partially from in-batch sampling and partially from the product pool, ensuring that all items in the product pool are exposed to the model during training. The loss is as follows:
@@ 3.2. Retrieval-GRPO
  • Candidate Selection: The dense retrieval model encodes queries and items into a unified semantic embedding space. Using the Approximate Nearest Neighbor (ANN) algorithm(Malkov and Yashunin, 2018), it retrieves the top-k items with the closest distances as candidate items.
@@ 3.2.1. Multi-Objective Reward
To satisfy users’ search experience, dense retrieval aims to prioritize items that are highly relevant to the user’s query and of superior quality. However, due to the prevalence of multi-way retrieval architectures in e-commerce search systems, we particularly emphasize incrementally exclusive relevant items—those that are semantically related to the query but not exact matches—that provide incremental value compared to results from other retrieval ways.
@@
where the $f_{\text{relevance}}(q,d)$ denotes the relevance score between query and item computed by TaoSR1, $g_{\text{quality}}(d)$ denotes the objective quality score of items derived from historical transaction data, and $h_{\text{exclusivity}}(q,d)$ measures exclusivity against other retrieval-ways, indicating whether the item provides incremental value by being retrieved through the current channel but not through alternative retrieval ways.
  • Relevance reward:The query-item relevance score is computed by TaoSR1(Dong et al., 2025), an advanced internal downstream relevance model deployed independently. TaoSR1 is a 42B-MoE LLM trained through multi-stage RL learning. We leverage it as the reward model in our reinforcement learning framework to provide real-time relevance scoring for query-item candidates generated by dense retrieval model.
@@ 3.2.2. Candidates Selection
In the Retrieval-GRPO framework, instead of offline hard negative samples construction, we treat the top-k items retrieved by the dense retrieval model as candidate predictions requiring correction. Leveraging the relevance reward model and multi-objective reward function described in Section 3.2.1, the rewards are computed in real time. Items within the top-k that receive low reward scores are penalized to dynamically adjust the model’s predictions.
@@ 3.2.3. GRPO loss
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\bigg\{\min\bigg[\frac{\pi_{\theta}(s(q,d_{i})|q,d_{i})}{\pi_{\theta_{\text{old}}}(s(q,d_{i})|q,d_{i})}\hat{A}_{i,t},\\ &\quad\quad\text{clip}\left(\frac{\pi_{\theta}(s(q,d_{i})|q,d_{i})}{\pi_{\theta_{\text{old}}}(s(q,d_{i})|q,d_{i})},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\bigg]-\beta\mathbb{D}_{\text{KL}}\big[\pi_{\theta}||\pi_{\text{ref}}\big]\bigg\}\Bigg]\end{split}$